Use the division algorithm please

3) Find the remainder when 1512+3 is divided by 13.
Answer #1

15 and 13 are coprime to each other.

And \phi (13)=13(1-\frac{1}{13})=12

So 15^{12}= 1 \ mod \ 13

Hence 15^{12}+3= 1+3 \ mod \ 13=4\ mod\ 13 Hence when 15^12 +3 is divided by 13 the remainder is 4